A carpenter purchased 70 ft of redwood and 90 ft of pine for a total cost of $325. A second purchase, at the same prices, included 100 ft of redwood and 50 ft of pine for a total cost of $370. Find the cost per foot of redwood and of pine. What is the price per foot of each?
Answer by richwmiller(17219)

70r+90p=325,
100r+50p=370
p = 6/5, r = 31/10
p=$1.20 r=$3.10